Astronomy Report – December 31st



Name: Gabriel Skowronek
Crew: 289
Date: 31 Dec 2023
Robotic Telescope Requested RCOS-16
Objects to be Imaged this Evening: SW Tauri
Images submitted with this report: None

Problems Encountered: None

Robotic Telescope Requested MDRS-WF
Objects to be Imaged this Evening: M1 (Crab Nebula), M42
Images submitted with this report: None

Problems Encountered:M1 has yet to image


Solar Features Observed: N/A

Images submitted with this report: N/A

Problems Encountered: N/A

Mid-Mission Research Report – december 30st

[category science-report]

Crew 289 – Deimos
Dec 24th, 2023 – Jan. 6th, 2023

A logo of a planet with planets and stars Description automatically generated

Crew Members:
Commander: Adriana Brown
Executive Officer and Crew Journalist: Sara Paule
Crew Geologist: Eshaana Aurora
Crew Engineer: Nathan Bitner
Health and Safety Officer and Crew Astronomer: Gabriel Skowronek
GreenHab Officer: Riya Raj

Crew Projects:

Title: Remote Station Monitoring
Author: Nathan Bitner
Objectives: Demonstrate usefulness of supplying on-site crew and mission control with real time habitat data
Current Status: We have created a prototype air quality monitor that successfully sends data to an online dashboard. Right now, this board stops functioning after any WIFI connection issues, requiring a manual reset, and quickly drains battery power. In the next two days we plan to fix these issues by putting the raspberry pi into idle mode when not in use and having the board automatically reset itself after any runtime errors. After our prototype is robust enough to function on its own we plan to create four more identical modules for other station rooms and then monitor airlock status on the same dashboard using magnetic reed switches.
EVAs: None

Title: Recording Dust Levels in the HAB
Author(s): Gabriel Skowronek
Objectives: Qualitatively determine the dust levels before and after cleaning procedures
Current Status: Several swab observations have been made throughout the Habitat to understand how much dust was present prior to arrival. In the lower deck of the Hab, swabbing was done underneath the First Aid station and on top of the shelf that holds the comms equipment. In the upper deck, swabbing was done on top of Pantry 3 and the top of the cabinet that has the comms equipment. After initially swabbing, the surfaces were cleaned and wiped down of any dust with a Clorox wet wipe to get a ‘clean’ baseline. In the coming days, swabbing will be done to track subsequent dust accumulation. Only one swab has been conducted thus far, 24 hours after surface cleaning and it was found that there was negligible dust accumulation with only a few dark fibers and fine light colored particles found on the cotton swabs. Swabbing will continue to see how much dust accumulates in the four chosen spots and will be compared against each other to see where dust settles the most.
EVAs: None

Title: Astronomy on Mars
Author(s): Gabriel Skowronek
Objectives: Determine the period of variation for SW Tauri, a Cepheid variable. Furthermore, it will be helpful to make impressive pictures of the Crab and Orion nebulae.
Current Status: Three photometry measurements have been taken of SW Tauri thus far. One has been taken per night but over the coming days, measurements will be taken twice per night at 4-hour intervals to acquire more data points for the light curve. The light curve will begin to take shape over the coming days as intensity measurements will begin to be plotted over time. The initial two measurements were slightly over exposed, so adjustments were made in the exposure time to get proper exposures. There have been some delays in imaging the Crab Nebula as the MDRS-WF seems to image several days after making the initial request. One set of images has been completed but came out to be over exposed, so a new request has been put in with adjusted exposure times and we are currently awaiting images to be taken. In the coming days, the Orion Nebula will also be imaged but further research will need to be done into the proper exposure times.
EVAs: None

Title: Comparison of Self-selection Traits versus Skill Utilization by Mars Colonists
Author(s): Sara Paule
Objectives: Examine the skills (e.g., flexibility, leadership, communication, problem-solving, domestic skills, etc.) most used by “colonists” in their day-to-day activities at the Mars Desert Research Station (MDRS) versus their ratings of importance pre- and post-mission.
Current Status: Pre-mission data has been downloaded from Qualtrics. The crew has completed 5 days-worth of daily surveys from Sol 1 (December 25, 2023) through today, Sol 6 (December 26, 2023). These were completed at the end of the mission day as planned. Paper survey results have been entered into an Excel database in preparation for post-mission data analysis. The crew will continue to complete daily surveys throughout mission and paper data will be converted to digital format. Post-mission surveys will be sent out after return to Earth.
Title: Establishing Best Practices in Mission Reporting from Prior Crew Reports
Author(s): Sara Paule
Objective 1: Establish the average word length of the various report styles.
Objective 2: Examine whether crew members are most often referenced by surname/family name, given name, or both.
Objective 3: Determine common subject matters within reports, such as references to meals, sleep, showering, etc.
Current Status: Pre-mission all the reports for the past calendar year were downloaded from the MDRS Reports webpage. A sample from each mission of the last year uploaded to the reports repository for both the Journalist Report and Sol Summaries have been randomly selected for analysis. Additionally, random samples of the Journalist Report and Sol Summaries were pulled from the emails of the prior crew (288) received pre-mission. In total, 15 crews were identified during that time period and 14 Journalist and 13 Sol reports were acquired using the aforementioned methods. Word length and character length have been calculated for each. Next steps will be to begin flexible qualitative coding to assess content.
EVAs: None

Title: Martian analog paleotemperature reconstruction
Author(s): Adriana Brown
Objectives: (1) Sample a measured section of sediments up the side of Hab Ridge and (2) identify the percent of carbonate present in sediments, (3) collect Pycnodonte fossils from the Tununk shale near Hab Ridge for carbonate clumped isotope analysis, (4) identify bentonite presence and frequency within the Tununk Shale, and (5) catalog and prepare samples for drilling.
Current Status: Progress on all objectives has been made. 36 fossils have been collected from Hab Ridge, cleaned, and catalogued. They are ready to return to Earth for geochemical analysis. The carbonate percent weight experiment is up and running with initial sediment samples from fossil collection sites dissolved in 0.1 M HCl and currently drying. 93 1.0 mL sediment samples from two measured sections of Hab Ridge have also been collected, and selected samples from these sections will be added to the percent carbonate experiment this upcoming week. These samples will be fully catalogued at MDRS and then will be ready to return to Earth for stable carbon isotope analysis at the University of Michigan. Lastly, several bentonite “swarm” locations have been noted, with beds documented at Barrainca Butte and sampled at Hab Ridge.
EVAs: Two EVAs to Hab Ridge to collect fossils and sample two measured sections, one EVA to Barrainca Butte to evaluate Pycnodonte presence and bentonite swarms.

Title: Mars Exploration by Origami Robot and Drone Scouting or Transportation
Author(s): Aditya Arjun Anibha
Objectives: Apply the concept of transformable origami robots that can exhibit multiple types of locomotion and test their ability to supplement exploration. Investigate feasibility of transporting robot using drone and scouting locations of interest prior to exploration.
Current Status: During EVA 3, the drone was tested for its carrying capacity within stability limits with cardboard box container carrying rocks with a suspended transparent fishing line harness to avoid sensor interference and keep payload at safer proximity than taping it onto the drone. It was able to carry up to 350 grams but wobbled due to swinging or when directly underneath the drone’s height sensor. The drone would therefore be better used to support the robot rather than carry it due to weight limits.
During EVA 7, the robot was tested for its ability to traverse mild rocky, uneven and sloped terrain. It was successfully made to travel at a speed of 0.3 m/s for 8 meters in its closed wheel configuration and 57 meters in its open wheel configuration, while supported by a tugging string to lighten its weight closer to Martian conditions. This was across uneven offroad terrain with varied distributions of rocks between 1 cm to 3 cm in diameter. It climbed three mounds with slope angles varying upto a maximum of 20 degrees. The robot’s total scale-measured mass on Earth is about 1.5 kg. Its effective scale-measured mass reduced to around 0.9 kg when vertically tugged or supported, which is higher than its expected scale-measured mass on Mars of 0.6 kg. Therefore, we can determine that it would operate freely without the tug on Mars.
The robot will now undergo maintenance before executing another EVA in Candor Chasma to test it with diverse terrain environments, where testing its jumping and peristaltic motion capabilities will be a priority. The drone will be used in place of a tugging string to support the robot’s motion and lighten its weight such that it experiences motion similar to that which it would on Mars
EVAs: EVA 3 for Drone Capacity Testing with Suspended Payload Harness and Drone Scouting of Candor Chasma conducted from Pooh’s Corner, EVA 7 for Origami Robot Exploration Testing at Cowboy Corner

Title: Miniaturized Martian Agrivoltaics
Author(s): Eshaana Aurora
Objectives: To comprehensively test the impact of solar and artificial irradiation on crop yields within an enclosed, module-like environment. To understand the feasibility of a miniaturized agrivoltaic farm within the MDRS Greenhab.
Current Status: The mini farm has been successfully assembled in a discreet corner of the Greenhab, addressing the low humidity issue caused by a few open spots with a makeshift solution—cling wrap placed on top of pots secured by rubber bands. Notably, the preliminary results indicate that the fully shaded Kale sapling has sprouted the first few shoots, while the non-shaded ones are yet to show signs of growth. Following successful troubleshooting and error management, the arduino and sensors, including temperature, IV Tracer, and Solar irradiation sensors, are now fully operational. The next phase involves attaching these sensors to the mini farm for real-time data collection. Over the coming days, a routine will be established, wherein data will be read and uploaded to Google Drive during the communication window, and plants will be watered and monitored for continued sprouting and growth.
EVAs: 0

Title: LiDAR Scanning of Terrain
Author(s): Riya Raj
Objectives: Will use phone apps to provide accessible scans of the terrain.
Current Status: I have been using Polycam on IOS to help me get more visuals on the structures that are near the MDRS campus. The most important thing to keep in mind is to include efficiency if you want to expand. For example, our recent crew EVAs were helpful in showing me large terrain and flat terrain that could potentially be used for solar farming or other habitats! Since MDRS is a growing program, we should look into things that will help with further research! I took an image scan and LiDAR scan of Hab Ridge and the MDRS campus. This lets us know what exists and what things could be improved for development. I have been loading 150+ scans during our COMMS windows to see how we can effectively image all structures. Most scans showcase layering, formations, and structure of the terrain!
LiDAR also helps with hazard assessments to scan what large rocks could pose a threat in areas of frequent visitors. As someone who is in Civil, I would like my work to reflect in what ways we can help the Earth and people. We need to learn to adapt and mitigate solutions to the current climate change issues. Expanding the campus with more renewable sources such as solar would be helpful!

Title: Radiation on Kale:
Author: Riya Raj
Objectives: Work on sustainable methods of growing veggies using simulations in hydroponics to provide fresh nutrients for the body.
Current Status: Hydroponics is a good example of controlled agricultural practices that can help increase plant growth time rates and heath. During my time at MDRS, I have been using 12-hour intervals with a control vs. variable experiment. The variable experiment includes the addition of H2O2 with the hydroponics module to compare the plant roots and leaves. I will be conducting image scans of the roots/leaves while monitoring the water pH/temp, surrounding temp/humidity, and plant cell structure with microscope views.
Hydrogen peroxide and radiation are distinct in their effects on plants. Radiation, such as ultraviolent (UV) radiation or ionizing radiation, can lead to the formation of reactive oxygen species (ROS) within plant cells. Similarly, hydrogen peroxide is a reactive oxygen species. When plants are exposed to stressors like radiation or hydrogen peroxide, they may experience oxidative stress, which can cause damage to cellular components such as proteins, lipids, and DNA.
Space exploration cannot happen without understanding what makes humans thrive. Learning from many agricultural practices here on earth has helped us look into sustainable ways to support human living for long term space missions. Bioregenerative life support systems are the way to include efficiency and optimization in human health. We should look further into how the extreme microgravity environment can impact plant growth. With bioregenerative life support systems, we can learn to regenerate water, oxygen, and food needed by the astronauts. I hope to help in the field of controlled agriculture practices to help countries in regions that are unable to produce proper crops.

Title: Indoor Air Quality
Author: Riya Raj
Objectives: Use sensors to monitor air quality
Update: This project will include basic air quality sensing with an EPA Dust Particle Sensor to analyze the intensity of PM2.5 within the air. The crew was successful in changing the air filter for the hab so I will be using the old filter to test the sensor. I am still waiting for other parts for the sensors to come in. I will also be building another sensor that will help detect the CO2, Ozone, TVOC, Relative Temp/Humidity, and PM 2.5 based on a specified time frame.
The importance of air quality is imperative for life support systems here on Earth, ISS, and future life support systems maybe on the Moon or Mars. Maintaining good indoor air quality is crucial for promoting a healthy, comfortable, and productive indoor environment, as well as preventing potential long-term health effects associated with exposure to indoor pollutants. An excess of compounds or particles in the air could cause dizziness, nausea, respiratory diseases, and many other dangerous health issues.
There are many countries suffering from the impact of climate change. Some are experiencing toxic smog due to inadequate transportation methods and other sources as we know of. Learning to properly ventilate areas and keep the air clean will not only keeps us healthy, but also make the Earth happy!

Operations Report – December 30st

SOL: 6
Name of person filing report: Nathan Bitner
Non-nominal systems: None
Notes on non-nominal systems: None
Spirit rover used: No
Hours: (Before EVA)
Beginning Charge: (Before EVA)
Ending Charge: (On return from EVA, before recharging)
Currently Charging:
Opportunity Rover used: Yes
Hours: 152.5
Beginning Charge: 100%
Ending Charge: 57%
Currently Charging: No
Curiosity Rover used: No
Hours: (Before EVA)
Beginning Charge: (Before EVA)
Ending Charge: (On return from EVA, before recharging)
Currently Charging:
Perseverance Rover used: Yes
Hours: 281.5
Beginning Charge: 100%
Ending Charge: 49%
Currently Charging: No
General notes on rovers: Rovers are functioning properly.
Summary of Hab operations:
Water (static tank): 387.6
Static tank pipe heater (on or off): on
Static tank heater (on or off): on
Toilet tank emptied: yes
Summary of internet: Internet is nominal
Summary of suits and radios: Suits and radios nominal.
Summary of Greenhab operations:
Heater: on
Supplemental light: off
Basil – 1 g
Summary of ScienceDome operations: Turned off small fan for hydroponics project.
Dual split: Heat or AC: On
Summary of RAM operations: (Tools used, work done):No work done. Used measuring tape and meter stick during EVAs for recording robot distance travelled and sediment height respectively. Both later returned.
Summary of any observatory issues: All systems nominal
Summary of health and safety issues: No health or safety issues
Questions, concerns and requests to Mission Support: None

Sol Summary – December 30st

Sol: 6
Summary Title: Sol Six: Robot Tricks and a Sediment Mix
Author’s Name: Adriana Brown
Mission Status: Nominal
Sol Activity Status:
Sol 6 began with the crew chipping away at the vat of oatmeal from yesterday morning’s breakfast. Soon after, the first EVA crew of the day, Adi (SCI), Sara (XO), and Nathan (ENG) suited up for a brisk 9:30 AM departure to Cowboy Corner. This EVA was an important one; the objective was to test the performance of E.L.F. As one may recall from earlier reports, E.L.F stands for Electronic LeapFrog. As such, the ability of E.L.F to maneuver over rough terrain was tested. We are happy to report that our little robot friend was able to navigate the rocky terrain, but future tests on jumping and crawling will be required.
45 minutes after the first EVA crew returned, the second EVA crew consisting of Adriana (CO), Eshaana (GEO), Gabe (HSO), and Riya (GHO) set out at noon to a familiar field site: Hab Ridge. At Hab Ridge, two sites were selected to bulk sample a stratigraphic section. The first section was at the base of Hab Ridge where the team, broken into two pairs, collected 45 samples over 75 feet of section. Moving farther up path 1103, the team found a second section to sample, where they collected 48 samples over a 93 foot section. Adriana will use these samples to test weight percent carbonate in the different strata and, back on Earth, run them for carbon isotopes as a proxy for oxygenation in seawater. Their sediment-sampling-efforts (a.k.a. a lot of crawling around in the dirt) were much appreciated by Adriana. Thanks, team!
The thoroughly-chilled EVA crew returned at 2:30 PM to the warm and cozy Hab where their crewmates had prepared some delicious black bean burger nuggets, expertly seasoned mashed potatoes (great job, Adi!), and leftover enchiladas. After a hearty lunch, much of the crew retired for short naps and to their various workstations. As we enjoy our sixth evening on the Red Planet, the crew is looking forward to Sara’s savory tomato basil pancakes and, most of all, sleeping in on our rest day tomorrow!
Look Ahead Plan: No EVAs tomorrow, Sunday will be used as a rest day
Anomalies in work: None
Weather: cold and cloudy with some sparse sun early in the morning
Crew Physical Status: Nominal
EVA: EVA #7 to test robotic capabilities and EVA #8 to sample two measured sections on Hab Ridge
Reports to be filed: Journalist Report, Greenhab Report, Operations Report, Sol Summary, Astronomy Report, Photos, EVA report, Mid-Mission Summary
Support Requested: None

Journalist Report – December 30st

Sara Paule, Crew 289 Journalist & Executive Officer
Sol 6
On Sunday, Mars astronauts rest, which means that Saturday is a full day when on a brief mission rotation. As such, all crew members went out on EVA today so we could conduct our seventh and eighth EVAs of mission. Ensuring we can complete our scientific objectives requires careful planning and management of personnel so that we can coordinate 1) EVAs, 2) research that needs to be conducted in the hab, and 3) normal day-to-day operations such as meals and cleaning. Eshaana a.k.a. Butter has taken point on EVA scheduling in cooperation with CO Adriana (Rocky) to manage geological sampling, LiDAR scanning, and testing of the Electronic Leaping Frog (E.L.F.) robot built by Scientist Michelin (Aditya) and the various steps along the way to get us to our final objectives.
After many days trouble-shooting E.L.F., Michelin finally got to take his baby out for a test run in the Martian terrain today. It was a moment of big anticipation for the whole crew. Shrub (Nathan) and Zinger (me) provided (moral) support. E.L.F. needed a little bit of an assist – simulating the reduced weight in Martian gravity – so Michelin and Zinger took turns feeding E.L.F. commands and gently tugging on his leash. E.L.F. performed successfully in loose rocky soil as well as on some larger rocks (~4 cm) with wheels initially in compact and then in expanded form. His true nemesis turned out to be a small desiccated plant, something he would not encounter on Mars proper, so it was a resoundingly successful first trial run, complete with astronauts crawling about to collect performance measurements. A LiDAR scan of the largest hill E.L.F. climbed will hopefully yield additional slope data at less cost to astronaut knees.
The afternoon EVA focused on geology and involved quite a bit more crawling, measuring, and also bonus digging. Rocky and Butter, assisted by Pots (Riya) and Terminator (Gabe) collected ~90 1mL tubes of sediment from Hab Ridge. Via the team’s close inspection and measurements, Rocky was able to gain a substantially improved understanding of the stratigraphy of the area. While out on EVA, Pots got to use a rock hammer for the first time ever and is now contemplating a change in career to become a geologist – not really but she did love smashing things with it. The crew at the top of the ridge also was able to capture some amazing views of the hab.
Other landmark moments in our scientific objectives include that Pots and Butter both have multipe sprouts now! The plant growth over the next week will help them measure the effects of radiation on growth and how minifarm shade affect growth of plants respectively. Pots also finished a lego and origami mock-up of her collapsible health station for use in the field. On the engineering front, Shrub was able to successfully troubleshoot the air sensor he built. While it is not yet working entirely as desired, he was able to collect CO2 data in the upper deck of the hab for the first time today. Terminator was able to determine exactly how to obtain intensity measurements for his variable star and is looking forward to graphing data points tomorrow to see how the light changes over multiple days. I also was able to make progress on both of my research projects today. The crew is using quite a few skills regularly each day – a mission like this requires a lot out of a person. Also, I now have Sol and Journalist samples from each crew over the past year to analyze and have word count data for those samples.
It is great to have accomplished so much in such a short period of time and today’s mid-mission report summary gave us a chance to reflect on what all we have accomplished to-date and to begin to plan for the next week (after a bit of a recharge tomorrow for rest day and New Year’s Eve).
It does feel like nearly everything on Mars is a chance to try something and see what works. Meals are certainly no exception. It is always hard to know exactly how much oatmeal to make and Nathan’s overestimate of yesterday yielded us breakfast for this morning, too. While EVA crew #2 was out playing in the sand, Shrub and Michelin teamed up to make some mashed potatoes and black bean burgers. They played with seasoning and both dishes turned out great. Meanwhile, I experimented with some more of the leftover oatmeal – adding chocolate chips, dehydrated strawberry dust from the bottom of a tin, brown sugar, and salt – to create an approximation of a cookie. While baking in the oven did yield a nice external crust, the inside remained a bit too moist to perhaps call the thing a “cookie” but there were no complaints about the flavor and none remaining post-lunch so the trial certainly was not a failure. For dinner tonight, inspired by the possibility of fresh basil from the GreenHab, I’ll be making savory tomato-basil pancakes. I can say, it smells delicious.

GreenHab Report – December 30st

GreenHab Officer: Riya Raj
Environmental Control: Heater on, fan off, door closed.
Average Temperatures:
Avg Temp – 9:39 am (68 °F)
Relative Humidity – 13 %
**After Watering:
Avg Temp – 9:43 am (66.2°F)
Relative Humidity – 66%
Used 3 gallon to water all the plants (includes misting amount)
Avg Temp – 4:12 pm (70.3°F)
Relative Humidity – 14%
**After Watering:
Avg Temp – 4:21 pm (68°F)
Relative Humidity – 55%
Used 2.8 gallons to water all the plants (includes more misting)
Hour of supplemental light: 10pm-2am (4 hours)
Daily water usage for crops: Sol 6 – 5.9 gallons
(includes 32 oz of sprayed mist at 10pm on 12/29)
Water in Blue Tank (200-gallon capacity): 174.85 gallons
Times of Watering for Crops: Sol 6 – 9:43 am & 4:21 pm
Changes to Crops: added large stakes to cucumber plants in the corner. Planted some new mint seeds and replanted the spinach seeds!
Sol 6:
The cling wrap works like a charm lolz. This is a great way to work around the dry humidity in the greenhab. The old leaves of the carrot plants are dying out while being replaced with new baby leaves and stems! The wrap on Eshaana’s plants are also working greatly because she also has a third sprout now! I moved over the wooden pallet to allow maximum sunlight on her teeny pots! My hydroponics experiemnet is also starting to have 10-15 more cracked seedlings and sprouts!!
I added some of the large stakes to help the cucumber plants in the giant blue tank pot stand up and eat some sunlight! The leaves were just growing on top of each other without any light shining on them. Now they are happy and tall! I then looked into the veggies and herbs that are already planted and I planted some mint! There is very little mint in the other pot so I used the rest of the seeds! Tomorrow, I will be planting something new as well…onions!!! I would like to see some more color in the greenhouse, so planting some flowers might be something to do in the future! Flowers are super pretty! ^-^ I will make an inventory or list of all the plants that have already been planted in the greenhab to prevent planting multiple veggies.
Sol 6 Harvest:
Basil: 1 gram (was a lot of leaves, I promise)
Support/supplies needed: None

EVA Report – December 30st

EVA Report #07
Author: Aditya Arjun Anibha
EVA Date: 30th December 2023
Start time: 0936
End time: 1116
Purpose of EVA:
Robot Functionality Field Testing at Cowboy Corner
Robot maneuverability will be tested using laptop controls near Cowboy Corner. Factors of consideration are roughness and steepness that it can handle, so overall efficiency of the robot movement will be considered, relative to conventional wheel rovers or human EVA.
The Extra-Vehicular Activity (EVA) commenced at 0936, led by Aditya (Crew Scientist), Sara (Crew Journalist) and Nathan (Crew Engineer). The crew embarked on their Martian exploration utilizing the Opportunity and Perseverance rovers, navigating from Cow Dung Road to the north until reaching Cowboy Corner. Opportunity and Perseverance initiated the mission with full charges – Opportunity at 100% and 152.5 hours, Perseverance at 100% and 281.8 hours.
After arriving at Cowboy Corner, the team determined a suitable travel path for the robot to cover a diverse set of terrain. Aditya set up the robot for testing, while Nathan would tug it via string and Sara would conduct video documentation. The experiment first started with the robot in its closed wheel configuration for flat terrain and was moved eastward for 8 meters. Next, the robot was changed to its open wheel configuration for rocky terrain and travelled eastward about 28 meters over a shallow mound, then northward 29 meters over a highly uneven and rocky mound. The crew then spotted a steeper rocky mound about 70 meters southward and tested the robot to climb over it. It successfully traversed these paths with varied distributions of rocks between 1 cm to 3 cm in diameter and slopes of upto 20 degrees from the horizontal. In both cases, the robot’s speed was about 0.3 m/s and the EVA crew followed it along its path. Small items were used as markers for the start and end points and the distance travelled between them was acquired using measuring tape.
With testing complete, the team travelled back to the rovers at Cowboy Corner and returned to the Habitat with Opportunity at 57% charge and 152.8 hours, and Perseverance at 49% charge and 281.8 hours. The reentry process for the HAB commenced at 1116, concluding their Martian exploration.
During testing, a few aliens were spotted on a distant mountain and three alien vehicles closer to North Ridge on the way back to the Hab.
This EVA involved meticulous planning, organization and time tracking to ensure the robotic exploration objectives were efficiently and successfully completed during the limited excursion time. It contributed many valuable insights into robot operation for exploration of varied flat, sloped, uneven and rocky terrains using variable wheel radii to overcome different degrees of rock size distributions. Overall, the EVA was executed with a resounding success with exploration, research and safety properly balanced and prioritized.
Destination: Cowboy Corner
Coordinates (use UTM WGS 84): N 518700, E 4253000
Participants: Aditya Arjun Anibha (Crew Scientist), Nathan Bitner (Crew Engineer), Sara Paule (Crew Journalist)
Road(s) and routes per MDRS Map: Drive north using Cow Dung Road 0110 up till Cowboy Corner
Mode of travel: Driving and Walking

Supplemental Operations Report – December 31st

Date: 12/31/2023
Name of person filing report: Sergii Iakymov
Reason for Report: Routine
Non-Nominal Systems: Intern trailer heater, Perseverance rear view mirror, Crew car

Power system:
Station Tank: 74%
Director Tank: 71%
Intern Tank: 81%
Generator Tank: 69%

Propane Readings:
Solar: Nominal.
SOC Last 24 hours: Max 100%; Min 70%; Avg 80.0%.
VDC Last 24 hours: Max 59.01V; Min 41.82; Avg 49.5V.
Generator run time: 4512.5 hours. Oil and air filter changed on 12/26/2023

Hab Static Tank – 387 gallons
GreenHab – 174 gallons
Outpost tank – 200 gallons
Science Dome – 0 gallons
Hab Toilet Tank emptied: Yes

Sojourner rover used: Yes, around campus.
Hours: 196.2
Beginning Charge: 100 %
Ending Charge: 100 %
Currently Charging: Yes
Notes on Rovers: Perseverance rear view mirror broken.

ATV: ATV’s Used: None. Nothing to report.

Hab Car used and why, where: To Hanksville for supplies. Need to be serviced.
Crew Car used and why, where: N/A
General notes and comments: Crew car towed to Loa for repairs.

Summary of Internet: All nominal.
EVA suits and radios: All nominal.
Campus wide inspection, if action taken, what and why: All nominal.
Summary of Hab Operations: All nominal.
Summary of GreenHab Operations: All nominal.
Summary of SciDome Operations: All nominal.
Summary of Observatories Operations: All nominal.
Summary of RAM Operations: All nominal.
Summary of Outpost Operations: Intern trailer heater isn’t functional and was dropped off for repairs on 12/22/2024
Summary of Health and Safety Issues: A lot of tourists trespassing starting December 22nd.

Copyright © The Mars Society. All rights reserved. | Main Site